# Characterizing Agent-Based Model Dynamics via $\varepsilon$-Machines and Kolmogorov-Style Complexity

### v2

@@ Abstract
We propose a two-level information-theoretic framework for characterizing the informational organization of Agent-Based Model (ABM) dynamics within the broader paradigm of Complex Adaptive Systems (CAS). At the macro level, a pooled $\varepsilon$-machine is reconstructed as a reference model summarizing the system-wide informational regime. At the micro level, $\varepsilon$-machines are reconstructed for each caregiver–elder dyad and variable, complemented by algorithm-agnostic Kolmogorov-style measures, including normalized LZ78 complexity and bits per symbol from lossless compression. The resulting feature set, $\{h_{\mu},C_{\mu},E,\mathrm{LZ78},\mathrm{bps}\}$, enables distributional analysis, stratified comparisons, and unsupervised clustering across agents and scenarios. Empirical results show that coupling $\varepsilon$-machines with compression diagnostics yields a coherent picture of where predictive information resides in the caregiving ABM. Global reconstructions provide a memoryless baseline (L=0L{=}0 under coarse symbolizations), whereas per-dyad models reveal localized structure, particularly for walkability under ordinal encodings (m=3m{=}3). Compression metrics corroborate these patterns: dictionary compressors agree on algorithmic redundancy, while normalized LZ78 captures statistical novelty. Socioeconomic variables display cross-sectional heterogeneity and near-memoryless dynamics, whereas spatial interaction induces bounded temporal memory and recurrent regimes. The framework thus distinguishes semantic organization (predictive causation and memory) from syntactic simplicity (description length) and clarifies how emergence manifests at different system layers. It is demonstrated on a caregiver–elder case study with dyad-level $\varepsilon$-machine reconstructions and compression-based diagnostics.
keywords:
Agent-Based Modeling , Complex Adaptive Systems , Computational Mechanics , $\varepsilon$-Machine , Statistical Complexity , Kolmogorov Complexity , Compression , Symbolic Dynamics , Entropy Rate , Information Theory
journal: Physica A
[…]
organization=Department of Informatics, Systems and Communication (DISCo), University of Milan-Bicocca,addressline=Viale Sarca 336,, city=Milan, country=Italy
@@ 2 Computational Mechanics for ABM: $\varepsilon$-Machines
In the context of ABMs, $\varepsilon$-machines provide a principled means to extract and quantify the system’s intrinsic computation—the transformation of past information into future behavior. Here, causal states may correspond to recurrent behavioral regimes, coordination cycles, or macro-level feedback patterns that emerge from agent interactions. Unlike scalar measures such as entropy or mutual information alone, the $\varepsilon$-machine reconstructs the minimal predictive model: it reveals the architecture of information storage and flow that governs the system’s evolution. This makes it particularly suited to studying adaptive or path-dependent ABM dynamics, where emergent regularities coexist with stochastic fluctuations. In contrast, Kolmogorov–Solomonoff–Chaitin complexity formalizes the descriptional compressibility of a sequence as the length of the shortest program that can reproduce it on a universal Turing machine [16]. Because true Kolmogorov complexity is non-computable, compression-based estimators such as Lempel–Ziv provide practical proxies [13].
The achieved compressed size reflects the regularity and redundancy of the sequence. A general compression-based complexity proxy is defined as:
@@
where $S_{\mathrm{comp}}$ is the compressed file size in bytes and $n$ is the number of symbols in the input sequence. The factor 8 converts bytes to bits, since $1~\text{byte}=8~\text{bits}$. This expression yields a value in bits per symbol, providing a dimensionless estimate of Kolmogorov complexity that is directly comparable to the entropy rate ($h_{\mu}$). Compression libraries such as lzma, bz2, and gzip report sizes in bytes, making this normalization necessary for consistency with information-theoretic measures. Lower values of $C_{\mathrm{comp}}$ indicate greater compressibility and structural regularity, while higher values reflect increased algorithmic novelty and unpredictability.
Both perspectives are complementary: the $\varepsilon$-machine captures semantic organization and predictive structure—the causal organization by which the system encodes, stores, and utilizes information to generate future states—whereas compression-based metrics capture syntactic simplicity, or the descriptive regularity of the system’s observable output.
@@ 3.1 Global $\varepsilon$-Machine as System-Level Reference
A single $\varepsilon$-machine is reconstructed from a pooled, symbolized time series (e.g., efforts) using quantile-based discretization. The optimal Markov order is selected by the Bayesian Information Criterion (BIC) up to $L_{\max}$, and histories are clustered by $L_{1}$ distance to obtain causal states. The resulting $(h_{\mu},C_{\mu},E)$ summarize the system-wide informational regime. When BIC selects $L=0$, the pooled process behaves approximately memoryless under the chosen resolution, providing a succinct contextual baseline for finer-grained analyses.
@@ 3.2 Per-Dyad, Multi-Variable $\varepsilon$-Machines and Algorithmic Proxies
Each dyad is treated as an independent stochastic process. For each variable—efforts, wkb, hrsncared (with a hurdle at zero), and binary overwhelmed—we reconstruct one $\varepsilon$-machine per dyad, obtaining $\{h_{\mu}^{(i,v)},C_{\mu}^{(i,v)},E^{(i,v)}\}$ and defining a complexity signature. These measures are complemented by algorithm-agnostic proxies: (i) normalized LZ78 complexity, and (ii) bits per symbol from lossless compression (LZMA, BZ2, GZIP). These proxies remain informative even when $\varepsilon$-machines collapse to trivial structures and provide clusterable feature sets for comparative and multivariate analysis.
@@ 3.3 Comparative Overview of LZMA, BZ2, and GZIP Compression Algorithms
The compression algorithms LZMA, BZ2, and GZIP all originate from the Lempel–Ziv family of lossless compression schemes, yet they differ substantially in their modeling depth, computational efficiency, and ability to detect long-range dependencies. When applied as proxies for Kolmogorov complexity, these differences translate into distinct sensitivities to structural regularity in symbolic sequences produced by agent-based models.
@@
where $S_{\mathrm{comp}}$ is the compressed file size in bytes and $n$ is the sequence length. This expression quantifies the bits per symbol required to encode the data. LZMA thus approximates long-range structural predictability, BZ2 captures intermediate regularities, and GZIP provides an estimate dominated by local redundancy. Using multiple codecs within the same analytical framework yields a multi-scale assessment of algorithmic regularity—analogous to evaluating informational structure at different temporal or spatial resolutions in entropy-based analyses.
@@ 3.4 Hierarchical Scopes and Visualization
4 System Level $\varepsilon$-machine Reconstruction
$h_{\mu}=H[X_{t}\mid X_{:t}],$ (8)
$C_{\mu}=H[\mathcal{S}],$ (9)
$E=I[X_{:t};\,X_{t:\infty}],$ (10)
Accurate estimation of these quantities, however, critically depends on how continuous simulation outputs are represented as discrete symbolic sequences. Symbolization—the transformation of numeric or multidimensional trajectories into finite alphabets—acts as the analytical lens through which temporal organization becomes observable. An appropriate scheme determines which variations are treated as equivalent and which are preserved as meaningful distinctions. Poorly chosen symbolizations can obscure dependencies by merging distinct states or, conversely, fabricate structure by discretizing noise into artificial categories. In agent-based simulations, this step is particularly delicate because variables differ in scale and generative mechanism: spatial indicators such as walkability exhibit smooth continuity and feedback, requiring ordinal or rank-based encodings that preserve relative change, whereas socioeconomic measures like caregiver effort or hours not cared vary primarily across agents and are better discretized using quantile or hurdle-mixture schemes. Symbolization thus defines the effective alphabet and the grain of observation, balancing detail and parsimony. Finer encodings may inflate entropy and complexity through over-partitioning, while coarser ones suppress correlations and underestimate structure. Beyond its technical role, symbolization is a modeling decision: it specifies the mapping between the simulated micro-dynamics and the informational architecture that $\varepsilon$-machines can reveal.
4.1 $\varepsilon$-Machine Reconstruction and Entropy Estimation Pipeline
4.1.1 $\varepsilon$-Machine Reconstruction
For each candidate order $L$, the empirical next-symbol probabilities are defined as:
$\hat{P}(a\mid h)=\frac{N(h,a)}{N(h)},\qquad N(h,a)=\big|\{\,t:s_{t-L:t-1}=h,\,s_{t}=a\,\}\big|,$ (11)
where $h$ denotes a history of length $L$. The log-likelihood and Bayesian Information Criterion are then given by:
$\log\mathcal{L}_{L}=\sum_{h}\sum_{a}N(h,a)\,\log\hat{P}(a\mid h),$ (12)
$\mathrm{BIC}(L)=-2\log\mathcal{L}_{L}+k_{L}\log N_{\mathrm{eff}},$ (13)
4.1.2 Predictive Distributions and History Clustering
For the selected order, predictive conditional distributions $P(a_{t+1}\mid a_{t-L:t})$ are estimated for all histories observed with sufficient frequency. Histories with statistically similar predictive distributions (as measured by their $L_{1}$ distance) are clustered to form equivalence classes representing causal states.
For each observed history $h$, we estimate:
$P(a\mid h)=\frac{N(h,a)}{N(h)},\qquad a\in\{0,\dots,A{-}1\},$ (14)
where $N(h,a)$ is the joint count of occurrences of history $h$ followed by symbol $a$, and $N(h)$ is the total count of $h$. Histories observed fewer than a threshold count are discarded to suppress noise. Two histories $h$ and $h^{\prime}$ belong to the same causal state if their $L_{1}$ distance satisfies
$P_{\mathrm{emit}}(a\mid s)=\frac{1}{|C_{s}|}\sum_{h\in C_{s}}P(a\mid h),$ (16)
defaulting to a uniform vector if no histories are present.
4.1.3 Unifilar Transitions and Stationary Distribution
A unifilar transition structure is a type of state–transition topology used in computational mechanics and predictive-state models (such as $\varepsilon$-machines), where each state–symbol pair uniquely determines the next state. A transition function $T(s,a)=s^{\prime}$ is constructed by majority rule over history shifts, producing a deterministic mapping between states and emitted symbols.
$h^{\prime}=\mathrm{shift}_{L}(h,a),$ (17)
and the next state $s^{\prime}$ is chosen as the most frequent state reached by these shifts. The corresponding state-transition matrix $M$ is defined as
$M_{ij}=\sum_{a}P_{\mathrm{emit}}(a\mid s_{i})\,\mathbf{1}\!\{T(s_{i},a)=s_{j}\},$ (18)
and it is normalized row-wise to ensure stochasticity. The stationary distribution $\pi$ satisfies
$\pi=\pi M,\qquad\pi^{(t+1)}=\pi^{(t)}M,\qquad\|\pi^{(t+1)}-\pi^{(t)}\|_{1}<\mathrm{tol}.$ (19)
4.1.4 Optional State Merging by Jensen–Shannon Divergence
To regularize over-fragmented models, states with similar emissions are merged when
$\mathrm{JS}\!\big(P_{\mathrm{emit}}(\cdot\mid s),\,P_{\mathrm{emit}}(\cdot\mid s^{\prime})\big)\leq\tau_{\mathrm{JS}},$ (20)
where $\mathrm{JS}(P,Q)$ is the Jensen–Shannon divergence and $\tau_{\mathrm{JS}}$ controls the merging threshold. The divergence is defined as
and after merging, the stationary distribution $\pi$ is recomputed.
4.1.5 Information-Theoretic Quantities
Entropy rate.
$h_{\mu}=\sum_{s}\pi_{s}\,H\!\big(P_{\mathrm{emit}}(\cdot\mid s)\big),\qquad H(p)=-\sum_{i}p_{i}\log_{2}p_{i},$ (22)
quantifying the average unpredictability of symbol generation.
Statistical complexity.
$C_{\mu}=H(\pi)=-\sum_{s}\pi_{s}\log_{2}\pi_{s},$ (23)
measuring the information stored in the stationary-state distribution.
Predictive information proxy.
Let
$P_{\mathrm{next}}=\sum_{s}\pi_{s}P_{\mathrm{emit}}(\cdot\mid s),\qquad H_{\mathrm{next}}=H(P_{\mathrm{next}}),$ (24)
then
$E_{\mathrm{proxy}}=\max\!\big(H_{\mathrm{next}}-h_{\mu},\,0\big),$ (25)
which approximates the system’s excess entropy, where $H_{\mathrm{next}}$ is the entropy of the one-step mixture distribution.
4.1.6 Finite-Block Entropy and Excess Entropy
An independent estimate of the entropy rate and excess entropy is computed via finite-block entropies up to order $L_{\max}$. For each block length $L$, the empirical entropy $H_{L}$ is estimated using a plug-in method with mild smoothing. Asymptotic values are extrapolated as
$h_{\mu}^{(\mathrm{block})}\approx H_{L}-H_{L-1},\qquad E^{(\mathrm{block})}\approx H_{L}-L\,h_{\mu}^{(\mathrm{block})}.$ (26)
$\hat{P}(b)=\frac{N(b)+1}{N+|\mathcal{B}^{L}|}.$ (27)
The block entropy is computed as
$H_{L}=-\sum_{b}\hat{P}(b)\log_{2}\hat{P}(b),$ (28)
and block-based estimates are refined by averaging the last three increments of $H_{L}-H_{L-1}$:
4.1.7 Visualization and Data Export
  1. (i)
    a time-series plot of the input symbols;
  2. (ii)
    a bar plot of stationary-state probabilities;
  3. (iii)
  4. (iv)
  5. (v)
    a table of BIC scores across candidate Markov orders.
The pipeline returns the chosen order $L^{\ast}$, the number of causal states $|S|$, the entropy rate $h_{\mu}$, statistical complexity $C_{\mu}$, predictive information proxy $E_{\mathrm{proxy}}$, block-based estimates $(h_{\mu}^{(\mathrm{block})},E^{(\mathrm{block})})$, and file paths for all generated visual and tabular artifacts.
4.2 Symbolization and Discretization Pipeline
Let the numeric series be denoted by
$\mathbf{x}=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}.$ (30)
A binning scheme with $k$ bins is defined by the ordered set of edges
$\mathbf{e}=(e_{0},\dots,e_{k}),\qquad e_{0}<\cdots<e_{k}.$ (31)
4.2.1 Discretization Methods
[…]
Fallback mechanisms ensure stability: Jenks $\rightarrow$ K-Means $\rightarrow$ quantiles for class breaks, and default clipping to small perturbations when numerical degeneracy occurs.
1. Quantile Bins.
$e_{j}=Q\!\left(\frac{j}{k}\right),\qquad j=0,\dots,k.$ (33)
2. Fixed or Policy-Fixed Edges.
3. Winsorized Equal-Width.
Using parameters $p_{\mathrm{low}}$ and $p_{\mathrm{high}}$, compute
$\ell=Q(p_{\mathrm{low}}),\qquad h=Q(p_{\mathrm{high}}),$ (34)
then clip as in Section 4.2.2. Edges are constructed as
$e_{j}=\ell+\frac{j}{k}(h-\ell),\qquad j=0,\dots,k.$ (35)
Alphabet size: $A=k$.
4. Jenks Natural Breaks.
Use edges $\mathbf{e}^{\star}$ obtained from the Jenks optimization in Section 4.2.2. Assign labels $\ell_{i}=\ell(x_{i};\mathbf{e}^{\star})$ and set alphabet size $A=k$.
5. Hurdle + Gaussian Mixture Model.
and shift labels by one to preserve the zero bin:
Alphabet size: $A=\hat{m}+1$.
6. Ordinal Patterns.
For an order parameter $m$, construct sliding windows
$\mathbf{w}_{t}=(x_{t-m+1},\dots,x_{t}).$ (40)
7. Lagged Vector Quantization.
For lag order $m$, build lagged vectors
$\mathbf{y}_{t}=(x_{t-m+1},\dots,x_{t})^{\top},\qquad t=m,\dots,n.$ (42)
Optionally standardize as
$\mathbf{z}_{t}=(\mathbf{y}_{t}-\hat{\mu})/\hat{\sigma}$, (43)
(elementwise). Apply $k_{\mathrm{VQ}}$-means clustering to minimize
$\sum_{t=m}^{n}\min_{j}\|\mathbf{z}_{t}-\mathbf{c}_{j}\|_{2}^{2}.$ (44)
Assign each vector to its nearest centroid:
$\ell_{t}=\arg\min_{j}\|\mathbf{z}_{t}-\mathbf{c}_{j}\|_{2}^{2}.$ (45)
4.2.2 Binning, Digitization, and Label Formation
Digitization.
Given edges $\mathbf{e}$, integer labels are computed as
Equal-Width Edges.
$e_{j}=\ell+\frac{j}{k}(h-\ell),\qquad j=0,\dots,k.$ (47)
Jenks Natural Breaks.
Jenks natural breaks [15] minimize the within-class sum of squares:
Winsorization.
$\ell=Q(p_{\text{low}}),\qquad h=Q(p_{\text{high}}),$ (50)
$x_{i}^{\text{win}}=\min\{\max\{x_{i},\ell\},\,h\}.$ (51)
The function returns $(\mathbf{x}^{\text{win}},\ell,h)$.

### v3

@@ Abstract
We present a two-level, information-theoretic approach for analyzing the temporal organization of outputs generated by agent-based models (ABMs), framed within the study of complex adaptive systems. At a system-wide level, a pooled $\varepsilon$-machine is reconstructed from simulated time series to provide a coarse baseline of overall predictability. At an agent level, $\varepsilon$-machines are estimated separately for each caregiver–elder dyad and model variable, and are paired with compression-based complexity measures derived from lossless coding and normalized LZ78 statistics.
These diagnostics support comparative and distributional analyses across agents, variables, and simulation scenarios. Applied to a simulated caregiving model, the results indicate that pooled dynamics are effectively memoryless under coarse representations, whereas dyad-level analyses recover localized temporal structure. In particular, spatial accessibility indicators exhibit bounded but nontrivial temporal organization when ordinal encodings are used, while socioeconomic variables primarily reflect cross-sectional heterogeneity with little short-range memory.
Compression-based measures are consistent with these findings: dictionary-based compressors yield concordant estimates of descriptive redundancy, while normalized LZ78 highlights differences in statistical novelty across variables. Taken together, the results illustrate how predictive structure and descriptive compressibility capture distinct aspects of simulated dynamics. The proposed approach is intended as a diagnostic framework for exploratory analysis of agent-based simulation outputs, rather than as a contribution to computational mechanics theory.
Keywords: Agent-Based Modeling; Complex Adaptive Systems; Computational Mechanics; $\varepsilon$-Machine; Statistical Complexity; Kolmogorov Complexity; Compression; Symbolic Dynamics; Entropy Rate; Information Theory.
@@ 1 Introduction: Complex Adaptive Systems and ABM
Scope and contributions
Despite the widespread use of agent-based models, the analysis of their outputs is still dominated by static summaries and low-order temporal diagnostics, such as averages, distributions, and autocorrelation functions. These tools capture variability but provide limited insight into how simulated agents store, transmit, or reuse information over time, particularly at the individual or dyadic level. As a result, there is a methodological gap between the rich temporal dynamics generated by ABMs and the comparatively coarse diagnostics typically used to interpret them.
This study is methodological in nature. It does not introduce new theoretical results in computational mechanics, nor does it aim to empirically validate agent-based models against observed social data. Instead, it addresses a practical and under-formalized problem in the analysis of agent-based simulations: how to diagnose and compare the temporal organization of simulated agent-level dynamics beyond standard summary statistics.
The paper makes three specific contributions. First, it operationalizes established tools from computational mechanics and information theory—namely $\varepsilon$-machine reconstruction and compression-based complexity measures—as diagnostics for time series generated by agent-based models. Second, it clarifies the distinct roles of predictive structure and descriptive compressibility by jointly analyzing $\varepsilon$-machine–based measures and algorithmic compression proxies, showing how these capture complementary aspects of simulated dynamics. Third, it demonstrates the approach on outputs from a caregiver–elder agent-based model, illustrating how spatially mediated variables exhibit bounded temporal organization at the agent level, while socioeconomic variables primarily reflect cross-sectional heterogeneity.
All results are conditional on the simulation design and representation choices adopted. The proposed framework is intended as an exploratory and comparative diagnostic tool for agent-based modeling studies, rather than as a contribution to the theory of complex adaptive systems or computational mechanics.
@@ 2 Computational Mechanics for ABM: $\varepsilon$-Machines
In the context of ABMs, $\varepsilon$-machines provide a principled means to extract and quantify the system’s intrinsic computation—the transformation of past information into future behavior. Here, causal states may correspond to recurrent behavioral regimes, coordination cycles, or macro-level feedback patterns that emerge from agent interactions. Unlike scalar measures such as entropy or mutual information alone, the $\varepsilon$-machine reconstructs the minimal predictive model: it reveals the architecture of information storage and flow that governs the system’s evolution. This makes it particularly suited to studying adaptive or path-dependent ABM dynamics, where emergent regularities coexist with stochastic fluctuations.
In contrast, Kolmogorov–Solomonoff–Chaitin complexity formalizes the descriptional compressibility of a sequence as the length of the shortest program that can reproduce it on a universal Turing machine [16]. Because true Kolmogorov complexity is non-computable, compression-based estimators such as Lempel–Ziv provide practical proxies [13]. The achieved compressed size reflects the regularity and redundancy of the sequence. A general compression-based complexity proxy is defined as
@@
where $S_{\mathrm{comp}}$ is the compressed file size in bytes and $n$ is the number of symbols in the input sequence. The factor 8 converts bytes to bits, since $1\penalty 10000\ \text{byte}=8\penalty 10000\ \text{bits}$. This expression yields a value in bits per symbol, providing a dimensionless estimate of algorithmic complexity that is directly comparable to the entropy rate $h_{\mu}$. Compression libraries such as lzma, bz2, and gzip report sizes in bytes, making this normalization necessary for consistency with information-theoretic measures. Lower values of $C_{\mathrm{comp}}$ indicate greater compressibility and structural regularity, while higher values reflect increased algorithmic novelty and unpredictability.
Throughout the remainder of the paper, any reference to reconstruction order or history length refers to the Markov order $K$. When an optimal order is selected via model selection criteria, it is denoted by $K^{\ast}$ and used consistently in all subsequent reconstruction steps unless stated otherwise.
Table 1: Notation used throughout the manuscript.
Symbol Meaning
$t$ Discrete simulation time index
$s_{t}$ Symbol observed at time $t$
$\mathcal{A}$ Symbol alphabet
$L$ Markov order (history length)
$L^{\ast}$ Selected Markov order (via BIC)
$L_{\max}$ Maximum Markov order considered
$s_{t-L:t-1}$ Length-$L$ symbol history ending at time t1t-1
$S$ Causal-state random variable
$\pi_{s}$ Stationary probability of causal state $s$
$P(\cdot\mid s)$ Emission distribution of state $s$
$\|\cdot\|_{1}$ $\ell_{1}$ (L1) norm
3 Model and data context
The analysis presented in this paper is based entirely on outputs generated by an agent-based simulation model. No empirical or observational social data are used. The purpose of the study is methodological: to demonstrate and evaluate diagnostic tools for characterizing the temporal organization of agent-based model (ABM) outputs under controlled simulation conditions.
Agent-based model overview
The underlying model represents a stylized caregiving system composed of caregiver–elder dyads. Each dyad consists of an older adult requiring assistance and an informal caregiver providing support. Agents interact within a simulated environment according to predefined behavioral rules that govern care provision, mobility, and access to services. Socioeconomic attributes (e.g., income or functional limitations) are fixed at initialization, while behavioral and accessibility-related variables evolve over simulation time.
The model is designed to generate heterogeneous agent-level trajectories rather than to reproduce or fit empirical observations. Simulation outputs are treated as realizations of a complex adaptive system, suitable for exploratory analysis of temporal structure and informational organization.
Simulation outputs and variables
Each simulation run produces time-stamped records for every caregiver–elder dyad. A dyad is the basic unit of analysis throughout the paper. For each dyad $i$ and simulation time step $t$, the model records several scalar variables. The present study focuses on the following outputs:
  • Caregiving effort (efforts): a continuous-valued indicator of the caregiving load or effort exerted by the caregiver at a given time step.
  • Walkability index (wkb): a scalar measure of spatial accessibility that summarizes the ease with which the dyad can reach essential services within the simulated environment.
  • Hours not cared (hrsncared): a non-negative variable representing unmet caregiving demand during a given time interval; this variable is zero-inflated by construction.
  • Overwhelmed status (overwhelmed): a binary indicator signaling whether the caregiver is classified as overloaded at a given time step according to model-specific thresholds.
All variables are generated endogenously by the simulation rules and are observed at a fixed temporal resolution.
Meaning of spatial structure
References to spatial or spatio-temporal structure in this paper refer specifically to the simulated physical environment in which agents move and interact. Spatial structure arises from the representation of locations, distances, and service accessibility within the model, and from the movement of agents across this environment over time. The walkability index captures the dynamic interaction between agent mobility and the spatial configuration of services. No abstract or metaphorical notion of space is implied.
Analytical scope
All analyses treat the recorded time series as simulation outputs to be diagnosed rather than as empirical signals to be explained or predicted. The goal is to characterize whether, and to what extent, different model variables exhibit temporal organization, memory, or redundancy under various symbolic representations. Any conclusions drawn are conditional on the simulation design and do not constitute claims about real-world caregiving systems.
@@ 4.1 Global $\varepsilon$-Machine as System-Level Reference
A single $\varepsilon$-machine is reconstructed from a pooled, symbolized time series (e.g., efforts) using quantile-based discretization. The optimal Markov order is selected by the Bayesian Information Criterion (BIC) up to $L_{\max}$, and histories are clustered using the $\ell_{1}$ norm to obtain causal states. The resulting $(h_{\mu},C_{\mu},E)$ summarize the system-wide informational regime. When BIC selects $L=0$, the pooled process behaves approximately memoryless under the chosen resolution, providing a succinct contextual baseline for finer-grained analyses. By pooled process we mean the concatenation of symbolized time series across all dyads along the temporal axis, after applying a common discretization scheme. This pooled sequence is used exclusively to construct a coarse, system-level reference model; it is not intended to represent agent-level dynamics, nor does it assume ergodicity or exchangeability across dyads.
@@ 4.2 Per-Dyad, Multi-Variable $\varepsilon$-Machines and Algorithmic Proxies
Each dyad is treated as an independent stochastic process. For each variable—efforts, wkb, hrsncared (with a hurdle at zero), and binary overwhelmed—we reconstruct one $\varepsilon$-machine per dyad, defining a complexity signature and obtaining $\{h_{\mu}^{(i,v)},C_{\mu}^{(i,v)},E^{(i,v)}\}$. These measures are complemented by algorithm-agnostic proxies: (i) normalized LZ78 complexity, and (ii) bits per symbol from lossless compression (LZMA, BZ2, GZIP). These proxies remain informative even when $\varepsilon$-machines collapse to trivial structures and provide clusterable feature sets for comparative and multivariate analysis.
Predictive versus descriptive complexity. $\varepsilon$-machines and compression-based measures quantify fundamentally different properties of symbolic processes. $\varepsilon$-machines characterize predictive equivalence classes of histories and therefore capture causal memory and forward-looking structure. In contrast, compression-based measures (LZ78, LZMA, BZ2, GZIP) quantify description length and redundancy, which need not imply predictability or causal organization. A sequence may be highly compressible yet weakly predictive, or vice versa. Their joint use is therefore diagnostic rather than redundant: agreement indicates robust structure, while divergence reveals representation- or scale-dependent effects.
Relation to Alternative Reconstruction Methods
Several alternative approaches exist for reconstructing predictive state models from symbolic time series. Notably, the Causal State Splitting Reconstruction (CSSR) algorithm, variable-length Markov models (VLMMs), and related subtree merging procedures provide principled means to infer state structure under assumptions similar to those underlying $\varepsilon$-machines.
These methods are widely used in computational mechanics and sequence modeling. However, they typically require larger sample sizes to stabilize state splitting, rely on multiple tuning parameters (e.g., significance thresholds or tree-pruning criteria), and produce reconstructions whose intermediate steps are less transparent when applied to short or heterogeneous sequences.
In the present study, the analytical objective is not to optimize predictive performance or to recover a unique generative model, but to diagnose the presence or absence of temporal organization across many short agent-level trajectories under controlled symbolic representations. For this reason, we adopt a direct history-clustering formulation of $\varepsilon$-machine reconstruction with explicit Markov-order selection and $\ell_{1}$-norm-based equivalence criteria.
This choice prioritizes interpretability, robustness under limited sample lengths, and comparability across dyads and variables. Alternative reconstruction algorithms are therefore excluded by design rather than by omission, and their use would not alter the qualitative role played here by $\varepsilon$-machines as diagnostic summaries of predictive structure.
@@ 4.3 Comparative Overview of LZMA, BZ2, and GZIP Compression Algorithms
The compression algorithms LZMA, BZ2, and GZIP all originate from the Lempel–Ziv family of lossless compression schemes, yet they differ substantially in their modeling depth, computational efficiency, and ability to detect long-range dependencies. When applied as proxies for algorithmic complexity, these differences translate into distinct sensitivities to structural regularity in symbolic sequences produced by agent-based models.
@@
where $S_{\mathrm{comp}}$ is the compressed file size in bytes and $n$ is the sequence length. This expression quantifies the bits per symbol required to encode the data. LZMA thus approximates long-range structural predictability, BZ2 captures intermediate regularities, and GZIP provides an estimate dominated by local redundancy. Using multiple codecs within the same analytical framework yields a multi-scale assessment of algorithmic regularity.
@@ 4.4 Hierarchical Scopes and Visualization
Throughout this section and the remainder of the paper, references to history length or reconstruction order refer to the Markov order $L$, while the notation $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm. For a given Markov order $L$, we denote by $s_{t-L:t-1}$ the length-$L$ history of symbols immediately preceding time $t$. When an optimal order is selected via BIC, it is denoted by $L^{\ast}$ and used consistently in subsequent analyses unless stated otherwise.